OpenStudy (anonymous):

Determine the solutions to the quadratic equation -10x2 + 30x - 20 = 0

OpenStudy (nurali):

-10x^2 + 30x - 20 = 0 (-10x^2 + 30x - 20)/(-10) = 0/(-10) x^2 - 3x + 2 = 0 (x - 2)(x - 1) = 0 x - 2 = 0 or x - 1 = 0 x = 2 or x = 1

OpenStudy (anonymous):

thanks (: